OpenStudy (anonymous):

Find the limit of the function by using direct substitution lim (2e^x sin x) x --> pi/2

OpenStudy (bahrom7893):

just plug in pi/2 into the equation.

OpenStudy (bahrom7893):

2e^(pi/2) * Sin(pi/2)

OpenStudy (anonymous):

such a simple question simply put x=pi/2

OpenStudy (anonymous):

what about e?

OpenStudy (bahrom7893):

use a calculator, e is a constant.

OpenStudy (bahrom7893):

about 2.71...

OpenStudy (anonymous):

I got 349 and that isn't an option on my choices :(

OpenStudy (anonymous):

Btw : \[\Huge \frac{\pi}2=\frac{180}2=90\]

OpenStudy (anonymous):

I have 0, 1, 2e^pi/2 and pi/2

OpenStudy (anonymous):

is pi/2 the answer or should I use 90 to solve the equation?

OpenStudy (anonymous):

because I used 3.14 as pi

OpenStudy (bahrom7893):

(2e^x sin x) x = 2e^(pi/2)*Sin(pi/2)*(pi/2) = 2e^(pi/2)*(pi/2)

OpenStudy (anonymous):

@Eyad

OpenStudy (bahrom7893):

= pi * e^(pi/2) that's the final answer

OpenStudy (anonymous):

thank you!!